# Emilia: An Extensive, Multilingual, and Diverse Speech Dataset for Large-Scale Speech Generation

### v1

@@ 2.1 Standardization
The raw speech data in the wild often vary in encoding formats, sampling rates, etc. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16-bit and adjust the target decibels relative to full scale to -20 dBFS. The required gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. These steps ensure a consistent data format for further processing.
@@ 2.2 Source Separation
The raw speech data in the wild often contain background music, which negatively impacts the performance of speech generation models [16, 17]. To address this, we use the source separation technique to extract clean human vocals. Specifically, we utilize the open-source library, Ultimate Vocal Remover, [^1] https://github.com/Anjok07/ultimatevocalremovergui and its pre-trained model, UVR-MDX-Net Inst 3.222https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models) This model achieves a high signal-to-distortion ratio of 11.15 for vocal separation on the Synth MVSep dataset. Using this model, we effectively separate human vocals from raw speech data for further processing.
@@ 2.5 ASR
The absence of reliable text transcriptions impedes the direct use of the Emilia dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [20], which is a state-of-the-art multilingual ASR model capable of speech translation and language identification. To further enhance efficiency, we leverage the faster-whisper [^5] https://github.com/SYSTRAN/faster-whisper backend using the CTranslate2666https://github.com/OpenNMT/CTranslate2 inference engine, which is up to four times faster than the official Whisper implementation for the same accuracy while using less memory. Additionally, we omit the original model’s VAD component by using the results from Sec. 2.4 to avoid redundant processing. We also develop batched inference for the faster-whisper backend to transcribe the speech data in parallel. These improvements, combined with the powerful Whisper model, allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.

### v2

@@ 2.1 Standardization
The raw speech data in the wild often vary in encoding formats, sampling rates, etc. To standardize the collected data, we convert all samples to WAV files, set them to a mono channel, and resample to 24 kHz. We set the sample width to 16-bit and adjust the target decibels relative to full scale to -20 dBFS. The actual gain is constrained within -3 to 3 dB to ensure appropriate volume without distortion. Finally, we normalize the waveform by dividing each sample by the maximum amplitude, ensuring values range between -1 and 1. These steps ensure a consistent data format for further processing.
@@ 2.2 Source Separation
The raw speech data in the wild often contain background music, which negatively impacts the performance of speech generation models [16, 17]. To address this, we use the source separation technique to extract clean human vocals. Specifically, we utilize the open-source library, Ultimate Vocal Remover, [^1] https://github.com/Anjok07/ultimatevocalremovergui and its pre-trained model, UVR-MDX-Net Inst 3 [18]. [^2] https://github.com/TRvlvr/model_repo/releases/tag/all_public_uvr_models) This model achieves a high signal-to-distortion ratio of 11.15 for vocal separation on the Synth MVSep dataset. Using this model, we effectively separate human vocals from raw speech data for further processing.
@@ 2.5 ASR
The absence of reliable text transcriptions impedes the direct use of the Emilia dataset for TTS. Therefore, we use ASR techniques to transcribe the segmented speech data. Considering the trade-off among speed, robustness, and accuracy, we employ the medium version of the Whisper model [21], which is a state-of-the-art multilingual ASR model capable of speech translation and language identification. To further enhance efficiency, we leverage the WhisperX [22], which builds on faster-whisper [^5] https://github.com/SYSTRAN/faster-whisper backend and the CTranslate2666https://github.com/OpenNMT/CTranslate2 inference engine, is up to four times faster than the official Whisper implementation for the same accuracy while using less memory. Additionally, we omit the original model’s VAD component by using the results from Sec. 2.4 to avoid redundant processing. We also develop batched inference for the faster-whisper backend to transcribe the speech data in parallel. These improvements, combined with the powerful Whisper model, allow our ASR step to achieve accurate text transcriptions for the speech data with high efficiency.